OpenStudy (anonymous):

Eliminate the parameter. x = 5 cos t, y = 5 sin t

OpenStudy (anonymous):

Same as before, what would you do this time?

jimthompson5910 (jim_thompson5910):

hint: square both sides and add the equations together

jimthompson5910 (jim_thompson5910):

why do this? because you can take advantage of the fact that sin^2(x) + cos^2(x) = 1

OpenStudy (anonymous):

so it would be x^2=25t?

jimthompson5910 (jim_thompson5910):

no not quite

OpenStudy (anonymous):

Think of your pathagoriean identities. And remember to factor out any constants...

jimthompson5910 (jim_thompson5910):

x = 5*cos(t) x^2 = (5*cos(t))^2 x^2 = (5^2)*(cos(t))^2 x^2 = 25cos^2(t)

jimthompson5910 (jim_thompson5910):

do the same for y

OpenStudy (anonymous):

so for y it would be y = 5*sin(t) y^2 = (5*sin(t))^2 y^2 = (5^2)*(sin(t))^2 y^2 = 25sin^2(t)

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

now add the two equations to get x^2 + y^2 = 25cos^2(t) + 25sin^2(t)

jimthompson5910 (jim_thompson5910):

see where this is going?

OpenStudy (anonymous):

x^2 + y^2 = 25[cos^2 (t) + sin^2(t) ] = 25

OpenStudy (anonymous):

so the answer would be what you put? @deepconcept

jimthompson5910 (jim_thompson5910):

that's not the final answer, but that's leading up to it

OpenStudy (anonymous):

final answer is x^2 + y^2 = 25 as cos^2(t) + sin^2 (t) = 1

OpenStudy (anonymous):

Thank you guys so much!!!

jimthompson5910 (jim_thompson5910):

you're welcome